# Dimensional reduction and adaptation-development-evolution relation in evolved biological systems

### v1

@@ Abstract
Life systems are complex and hierarchical, with diverse components at different scales, yet they sustain themselves, grow, and evolve over time. How can a theory of such complex biological states be developed? Here we note that for a hierarchical biological system to be robust, it must achieve consistency between micro-scale (e.g. molecular) and macro-scale (e.g. cellular) phenomena. This allows for a universal theory of adaptive change in cells based on biological robustness and consistency between cellular growth and molecular replication. Here, we show how adaptive changes in high-dimensional phenotypes (biological states) are constrained to low-dimensional space, leading to the derivation of a macroscopic law for cellular states. The theory is then extended to evolution, leading to proportionality between evolutionary and environmental responses, as well as proportionality between phenotypic variances due to noise and due to genetic changes. The universality of the results across several models and experiments is demonstrated. Then, by further extending the theory of evolutionary dimensional reduction to multicellular systems, the relationship between multicellular development and evolution, in particular the developmental hourglass, is demonstrated. Finally, the possibility of collapse of dimensional reduction under nutrient limitation is discussed.
@@ 1 Introduction
Biological systems generally consist of a huge number of components, and are hierarchical: cells consist of diverse molecules, organisms consist of diverse cells, and ecosystems consist of diverse species[1]. In order to have robustness in a biological system, there must be some consistency across levels; for example, in order for a cell to grow, almost all the components in it must be produced in proportion to their abundances; otherwise the cellular state is not robust. In the field of universal biology that we intend to establish, we seek to uncover universal laws in adaptation, development, and evolution by noting robustness as well as plasticity to adapt to environmental changes. Here, we briefly review recent advances in the field, focusing on the evolutionary dimensional reduction of high-dimensional biological states, as well as the correlation between the changes in evolution and adaptation or development. Last, possible violation of the dimension reduction is discussed in relationship with the transition to dormant (non-growing) state of microorganisms.
2 Dimension reduction as a consequence of robustness
@@
Recalling the diversity of components in a biological system, its state is high-dimensional: in a cell there are typically thousands of components. Then, the cellular state to be described by the concentration of each component requires a quite high-dimensional space; for a bacterial cell, about a few thousand protein or mRNA species (besides more metabolites), and the dimension of the state space exceeds at least a few thousand. Recent experimental and simulation studies, however, suggest that a drastic reduction in the dimensionality of the state change (phenotypes) in response to environmental changes has taken place[2-7].
@@
First, it is noted that the steady state imposes some constraint on the phenotypic changes[2], but it alone cannot explain the experimentally observed dimensional reduction in the changes of many components across different environmental changes.
@@
Then, we simulated a cell model consisting of a large number of components with catalytic-reaction network [3], in which phenotypes are generated as a result of intracellular reaction dynamics, where the abundances of each component change with catalytic reactions by taking nutrient chemicals from environment, leading to the cellular growth. When we adopted arbitrarily chosen reaction dynamics in the model, such reduction was not observed, even when the steady-growth condition of the cells was satisfied. In contrast, when these model cells were evolved to achieve a higher cell growth under a given environmental condition, it is found that the phenotypic changes of the cells upon a variety of environmental perturbations are constrained within a low-dimensional space.
@@
Accordingly, the fitted state (attractor) should be attracted from most directions in the high-dimensional state space. In terms of dynamical systems, this implies that there is a strong attraction from most directions to the stationary state corresponding to the fitted state. However, if this strong attraction occurred from all directions, evolution to increase the fitness would be difficult because the state would not be easily changed by genetic mutations, which slightly alter the dynamical system (the flow in the state space). Then, along the direction of evolution, the phenotypes should be feasible to be changed. Accordingly, dominant changes in phenotypes (e.g., concentration of proteins) will be constrained along a one-dimensional curve (or few-dimensional space) in which evolution has progressed and will pregress (see Fig. 1).
In fact, we confirmed the above behavior in the evolution of a cell model with catalytic reaction dynamics by considering the relaxation dynamics to the steady state. By linearizing the dynamics around the steady state, one (or few) eigenvalues for the relaxation matrix (Jacobi matrix) is close to zero, as distinct from other eigenvalues which are much more negative [6].
Besides the cell model with catalytic reaction network, the evolutionary dimensional reduction has also been confirmed in the evolution of gene-regulatory network models with mutual activation or suppression of gene expressions[8], where fitness is determined by the expression levels of given target genes, depending on environmental conditions. In both models, the dynamics involve many variables that provide the high-dimensional phenotypic dynamics, while the fitness for evolution (e.g., the growth rate of a cell) is determined from these phenotypes. In these numerical studies, it was found that after evolution, changes in most variables induced by a variety of different environmental perturbations are highly constrained along a one-dimensional manifold.
Here, with the dimensional reduction, the relaxation dynamics towards the steady state consists of one (or few) slow modes. The evolution of such a slow mode is reasonable by the following argument: Consider a complex system with many interacting elements with similar time scales. It will be more difficult to evolve such system, because the random mutations to each element (gene) would mutually interfere. Then, directed phenotypic changes by genetic changes would be harder, as the proverb says too many cooks spoil the broth. In contrast, if few slow modes are separated, they will work for controlling others. In other words, the separation of a few slow modes, which is a result of evolution, also facilitates phenotypic evolution.
@@ 3 Relationship between environmental and evolutionary responses
Since the one-dimensional curve for the phenotypic changes in response to environmental changes follows the curve along which the evolution has progressed, the responses due to environmental and evolutionary (genetic) changes will occur along the same low-dimensional constraint (see Fig. 1b). Thus, phenotypic changes due to evolution are highly correlated with those due to environmental or stochastic perturbations, as confirmed by the simulation of the cell models.
Here let us denote the change in phenotype (e.g., the log concentration of each component) due to environmental responses as , and that due to genetic evolution as . Since all phenotype changes are then constrained along the common one-dimensional curve, and will follow, where is the change in the principal mode and is the angle between and . Then, will be independent over all components . ( is the case in which the principal axis is linear. If it is curved, one can use a function that represents the projection of with the principal axis, and the same result is obtained).
@@
Since , this equation implies that the change in phenotype (e.g., logarithm of the concentrations of a protein or an mRNA species) induced by the stress is reduced as evolution progresses under this environmental stress, as . Thus, the long-term evolutionary response is to counteract the environmentally induced change. This homeostasis by adaptation is akin to the Le Chatelier principle in thermodynamics[4, 7]. In fact, eq. (1), together with the Le Chatelier-type responses, have been confirmed both in the simulation of cell models and in the evolution experiment of bacteria[4, 5].
@@ 4 Proportionality between the phenotypic variances by noise and by mutation
In considering the phenotypic variance, however, there is another type of fluctuations which are due to genetic changes (see FIg.2). Indeed, in standard population genetics, the genetic variance , which is the variance of the phenotype due to the genetic distribution, is known to be proportional to the rate of evolution[17, 18]. Accordingly, the proportionality between the phenotypic variances due to noise and due to genetic change is proposed, which is confirmed in the numerical simulation of the cell model over the course of evolution[19, 20].
Now, recall the proportionality of phenotypic responses by external environment and by genetic variation across components. We can then expect the proportionality of two phenotypic variances of by noise and by mutation, across components . This can be shown as follows: We denote the variance of the former by and that of the latter by . As shown in Fig. 1b, changes due to noise and genetic variation progress along the common principal axis. Thus, phenotypic variances due to the former are given by and those due to the latter by , where and are the average over the distribution of phenotypes induced by noise and mutations, respectively. Recall that the relaxation of phenotypic changes is much slower along the principal axis in Fig.1, phenotypic fluctuations due to noise are nearly confined to this axis. Thus, the variance of each component is given by the mean of the variance in the variable (along the principal axis) as (due to noise) and (due to mutation). Hence . and we obtain
@@
This proportionality over most of the components was confirmed by using the aforementioned cell model with catalytic reaction network [4], as well as in the evolution of the gene regulatory network model[16, 21], after the evolution has proceeded to achieve robustness,
Note that this result introduces the possibility of predicting which component is likely to evolve: changes due to noise itself are not inherited, but genetic variation is naturally transmitted to offspring, and the resulting phenotypic variance is proportional to the evolutionary rate of that component [17, 18]. However, since is proportional to , the rate of evolution of is proportional to . In other words, as the variance of the component (trait) due to noise or environmental change is larger in prior to the genetic change, the component (trait) is more feasible to evolve. In other words, the direction in which it is likely to evolve in phenotypic space can be predicted in prior to the genetic evolution[7].
@@ 5 Summary of evolutionary dimensional reduction and - law
In summary, we have demonstrated the evolutionary dimensional reduction and 2-by-2 global proportionality of phenotypic changes between responses and fluctuations, and between perturbations due to environmental (noise) and genetic changes, as schematically shown in Fig. 3.
We expect that this evolutionary dimension reduction and the resulting constraint on phenotypic evolution will be universal in evolved biological systems. In the theoretical argument and cell models, the prerequisites for the emergence of dimension reduction are only the following two: (i) phenotypes with higher fitness are shaped by complex high-dimensional dynamical systems, which involve stochasticity and (ii) the fraction of such fit states is rare in the state space and in the genetic-rule space. These two features are necessary for our theoretical argument, while most biological systems satisfy them. Intuitively, a state that satisfies macro-micro consistency must be rare, and once it is achieved, its conservation is favored, implying the requirement of robustness.
@@ 6 Universality
Universality of evolutionary dimensional reduction, correlation between environmental and genetic changes, and the proportionality between genetic variance () and isogenic phenotypic variation () has recently been examined in theoretical models and experiments. Here we briefly review these.
@@
The dimensional reduction by evolution has also been observed in several models. First, in reaction network models, phenotypic change is mainly constrained along a one-dimensional manifold across different types of fitness for selection (e.g., by selecting the concentration of a component). Second, even when the environmental conditions (e.g., the concentration of external nutrients) are not fixed but fluctuate over generations, the constraint within a one-dimensional manifold is observed [6]. Third, by using the gene expression dynamics model and by imposing a fitness condition depending on the expression pattern of target genes, the dimensional reduction and the relationship between variances are confirmed[8].
@@
Through a combination of data analysis, elastic network models, and theoretical analysis, we have confirmed the relationship between and , as well as the dimensional reduction of protein structure. Changes in estimated structure due to conformation dynamics under noise and genetic mutations are constrained to a common low-dimensional space, typically about five dimensions, implying a strong reduction from the hundreds of residues that give rise to the high-dimensional state space [22]. A theoretical explanation is proposed there, while the consequence of few slow modes has been discussed earlier [23], and some model studies also suggest such dimensional reduction [24].
@@
In relationship, we have also studied the evolving spin-Hamiltonian model. Here the spin variable takes up (+1) or down (-1) values, whose dynamics is governed by the interaction with other spins, with the energy , for spin and . (If is positive (negative), the two spins tend to be aligned (opposite)). By formulating the above interaction using a Hamiltonian () for energy, with the distribution under temperature (), where fitness is determined by a particular configuration of target spins and the matrix undergoes evolution by mutation and selection [25]. Under an intermediate temperature range, the evolutionary dimensional reduction and the relationship are then numerically confirmed [26, 27].
@@
(III) Experimental confirmations: Dimensional reduction was first observed in the transcriptome analysis of bacteria E coli, under different types of stress[2]. Support for equation (1) was obtained by the bacterial evolution under ethanol stress[4, 5]. In addition, transcriptome analysis oacross thousands of mRNA expression data showed that changes in bacterial responses under different antibiotics could be predicted by about seven principal components [9]. The cross-fitness after evolution under different types of antibiotics is shown to be correlated with the response in prior to evolution, as predicted by the theory[8]. Further support for such dimensional reduction comes from Raman analysis of cellular components [28].
@@
The proportionality between genetic and isogenic phenotypic variances across genes is also consistent with measurement of yeast gene expressions. Landry et al.[29] made global measurements of isogenic expression variation and mutation-induced variance. Although they used βmutational varianceβ (i.e., the rate of diffusive spread of each gene expression level upon the introduction of mutations) instead of , a high correlation between the isogenic phenotypic fluctuation () and the mutational variance was observed across all genes in yeast (see also [30]). More recently, the quantitative analysis of the gene expression pattern in Japanese medaka (fish) [31] has suggested proportionality.
@@
The application of the proportionality between variances theory is not limited to gene expressions, but can be applied to any other phenotypic trait. The correlation between the isogenic variance and the evolution speed was confirmed in the evolution experiment of RNA structure[32]. In a series of experiments, Stearns et al. [33] measured the isogenic variance of five life-history traits in Drosophila melanogaster, while also measured the genetic variance in the same five traits between different genetic lines. Interestingly, they observed proportionality between the isogenic variance of each trait and its genetic variance. Further support for the - proportionality can be found in the in the wing morphology of Drosophila (fly) [34, 35].
(IV) Related topics:
@@
The dimensional reduction has also been observed in the development of C. elegans[36] and in ecological evolution[37]. One possible link to the dimensional reduction will be the sloppy parameter hypothesis of Sethna [38], which suggests that many parameters used in biological models are irrelevant.
@@
Besides the issue in biophysics, dimension reduction is attracting much attention in learning in the brain [39] and neural networks [40], as well as in machine learning. In deep learning, the possible network configurations that can support a desired output are quite redundant, and reduction to the appropriate output space is achieved. This is similar to our case. The genetic space (e.g., the network configuration) is high-dimensional, so there can be many ways to achieve the appropriate phenotype. Even though the final selected states are constrained in low-dimensional space, the use of high-dimensional space is essential for both machine learning and evolution.

### v2

@@ Abstract
Living systems are complex and hierarchical, with diverse components at different scales, yet they sustain themselves, grow, and evolve over time. How can a theory of such complex biological states be developed? Here we note that for a hierarchical biological system to be robust, it must achieve consistency between micro-scale (e.g. molecular) and macro-scale (e.g. cellular) phenomena. This allows for a universal theory of adaptive change in cells based on biological robustness and consistency between cellular growth and molecular replication. Here, we show how adaptive changes in high-dimensional phenotypes (biological states) are constrained to low-dimensional space, leading to the derivation of a macroscopic law for cellular states. The theory is then extended to evolution, leading to proportionality between evolutionary and environmental responses, as well as proportionality between phenotypic variances due to noise and due to genetic changes. The universality of the results across several models and experiments is demonstrated. Then, by further extending the theory of evolutionary dimensional reduction to multicellular systems, the relationship between multicellular development and evolution, in particular, the developmental hourglass, is demonstrated. Finally, the possibility of collapse of dimensional reduction under nutrient limitation is discussed.
@@ 1 Introduction
Biological systems generally consist of a huge number of components, and are hierarchical: cells consist of diverse molecules, organisms consist of diverse cells, and ecosystems consist of diverse species[1]. In order to have robustness in a biological system, there must be some consistency across levels; for example, in order for a cell to grow, almost all the components in it must be produced in proportion to their abundances; otherwise the cellular state is not robust. In other words, the system has to satisfy macro-micro consistency; the cell growth at a macroscopic level need to keep consistency with the replication rate of each of constituent molecules at a microscopic level, when a cell achieves a state with robust growth. Likewise, for a multicellular organism, multicellular growth at an organism level needs to keep some consistency with dynamics of each cell. In the field of universal biology that we intend to establish, we seek to uncover universal laws in adaptation, development, and evolution by noting robustness as well as plasticity to adapt to environmental changes. Here, we briefly review recent advances in the field, focusing on the evolutionary dimensional reduction of high-dimensional biological states, as well as the correlation between the changes in evolution and adaptation or development. Last, possible violation of the dimension reduction is discussed in relationship with the transition to dormant (non-growing) state of microorganisms.
2 Dimensional reduction as a consequence of robustness
@@
Recalling the diversity of components in a biological system, its state is high-dimensional: In a cell, there are typically thousands of components. Then, the cellular state to be described by the concentration of each component requires a quite high-dimensional space; for a bacterial cell, about a few thousand protein or mRNA species (besides more metabolites), and the dimension of the state space exceeds at least a few thousand. Recent experimental and simulation studies, however, suggest that a drastic reduction in the dimensionality of the state change (phenotypes) in response to environmental changes has taken place[2-7], whereas the dimensional reduction of evolved phenotypes has been observed extensively in a variety of contents, as will be discussed in Β§6.
@@
First, it is noted that the steady state imposes a constraint on the phenotypic changes so that the cell growth rate and replication rates of molecules are balanced[2], but it alone cannot explain the experimentally observed dimensional reduction in the changes of many components across different environmental conditions.
@@
Then, we simulated a cell model consisting of a large number of components with catalytic-reaction network [3], in which phenotypes are generated as a result of intracellular reaction dynamics, where the abundances of each component change with catalytic reactions by taking up nutrient chemicals from the environment, leading to cellular growth. When we introduced arbitrarily chosen reaction dynamics into the model, no such reduction was observed, even when the steady growth condition of the cells was satisfied. In contrast, when these model cells were evolved to achieve higher cell growth under a given environmental condition, it is found that the phenotypic changes of the cells upon a variety of environmental perturbations are constrained within a low-dimensional space.
@@
Accordingly, in the dynamics for the phenotype, there should be a strong attraction to the fitted state (attractor) from most directions in the high-dimensional state space. In terms of dynamical systems, this implies that there is strong attraction from most directions to this fitted attractor. However, if this strong attraction occurred from all directions, evolution to increase fitness would be difficult because the state would not be easily changed by genetic mutations that slightly alter the dynamical system (the flow in the state space). Then, along the direction of evolution, the phenotypes should be feasible to change. Accordingly, dominant changes in phenotypes (e.g., concentration of proteins) will be constrained along a one-dimensional curve (or within a few-dimensional space) in which evolution has progressed and will progress (see Fig. 1).
In fact, we confirmed the above behavior, by the evolution simulation of a cell model with catalytic reaction dynamics mentioned above. By linearizing the dynamics around the steady state, one (or few) eigenvalues for the relaxation matrix (Jacobi matrix) is close to zero, as distinct from other eigenvalues which are much more negative. The evolutionary change has progressed along the eigenvector for the eigenvalue close to zero[3, 6].
In addition to the cell model with catalytic reaction network, the dimensional reduction has also been confirmed in the evolution of gene-regulatory network models with mutual activation or suppression of gene expression levels[8]. In the gene-expression model, fitness is determined by the expression levels of given target genes, depending on environmental conditions. In these models also, the dynamics involve many variables that provide the high-dimensional phenotypic dynamics, from which the fitness for evolution is determined. In the numerical studies, it was found that after evolution, changes in most variables induced by a variety of different environmental perturbations are highly constrained along a one-dimensional manifold.
Here, with the dimensional reduction, the relaxation dynamics towards the steady state involves one (or few) slow modes. The evolution of such a slow mode is reasonable by the following argument: Consider a complex system with many interacting elements that have similar time scales. It will be more difficult to evolve such a system because the random mutations to each element (gene) would interfere with each other and cancel each other out. Then, directed phenotypic changes by genetic changes would be more difficult, as the saying goes too many cooks spoil the broth. On the other hand, if a few slow modes are segregated, they will work to control others. Then mutations in the genes for the slow modes will strongly affect fitness. In other words, the separation of a few slow modes, which is a result of evolution, also facilitates phenotypic evolution.
@@ 3 Relationship between environmental and evolutionary responses
Since the one-dimensional curve for the phenotypic changes in response to environmental changes follows the curve along which the evolution has progressed, the responses due to environmental and evolutionary (genetic) changes will occur along the same low-dimensional manifold (see Fig. 1b). Thus, phenotypic changes caused by mutation are highly correlated with those caused by environmental or stochastic perturbations. This correlation was confirmed by the simulation of the cell models.
Here let us denote the change in phenotype (e.g., the log concentration of each component) due to environmental responses as , and that due to genetic evolution as . Since all phenotype changes are then constrained along the common one-dimensional curve, and will follow, where is the change in the principal mode and is the angle between and . Then, will be independent of each component . ( is the case in which the principal axis is linear. If it is curved, one can use a function that represents the projection of with the principal axis, and the same result is obtained).
@@
Since , this equation implies that the stress-induced change in phenotype (e.g., the logarithm of the concentration of a protein or mRNA species) is reduced as evolution proceeds under that environmental stress, since . Thus, the long-term evolutionary response is to counteract the environmentally induced change of . This homeostasis by evolutionary adaptation is akin to the Le Chatelier principle in thermodynamics[4, 7]. In fact, eq. (1), together with the Le Chatelier-type responses, has been confirmed both in the simulation of cell models and in the evolution experiment of bacteria[4, 5]. This Le Chatelier-type response can be understood by assuming that the fitness (growth rate) is described as a smooth, potential function of reciprocal variables representing environmental and genetic changes[8, 12].
@@ 4 Proportionality between the phenotypic variances by noise and by mutation
In considering the phenotypic variance, however, there also exists another source that causes the phenotypic fluctuations. It is the change in genes that govern the gene expression dynamics(see Fig.2). In fact, in the standard population genetics, the genetic variance , which is the variance of the phenotype due to the genetic distribution, is known to be proportional to the rate of evolution[18, 19]. Accordingly, the proportionality between the phenotypic variances due to noise and due to genetic change is proposed, which is confirmed in the numerical simulation of the cell model over the course of evolution[20, 21].
Now, recall the proportionality of phenotypic responses by external environmental changes and by genetic variation across all components. We can then expect the proportionality of two phenotypic variances of by noise and by mutation, across all components . This can also be shown as follows: Let us denote the variance of the former by and that of the latter by . As shown in Fig. 1b, changes due to noise and genetic variation are constraiend along the common principal axis. Thus, phenotypic variances due to the former are given by and those due to the latter by , where and are the average over the distribution of phenotypes induced by noise and mutations, respectively. Recall that the relaxation of phenotypic changes is much slower along the principal axis in Fig.1, phenotypic fluctuations due to noise are nearly confined to this axis. Thus, the variance of each component is given by the mean of the variance in the variable (along the principal axis) as (due to noise) and (due to mutation). Hence, follows. Accordingly
@@
This proportionality over most of the components was confirmed by using the aforementioned cell model with catalytic reaction network [4], as well as in the evolution of the gene regulatory network model[17, 22], after the evolution had proceeded to achieve robustness.
Note that this result introduces the possibility of predicting which component is likely to evolve: changes due to noise are not inherited, but genetic variation is transmitted to offspring, leading to the evolutionary change. Accordingly, the resulting phenotypic variance is proportional to the rate of evolution of that component, known as Fisherβs fundamental theorem of natural selection [18, 19]. Now that we have shown that is proportional to , the rate of evolution of is proportional to . In other words, the greater the variance of component (trait) due to noise or environmental change (before genetic change), the more likely component (trait) is to evolve. In other words, the direction in which it is likely to evolve in phenotypic space can be predicted in advance of genetic evolution[7].
@@ 5 Summary of evolutionary dimensional reduction and - law
In summary, we have demonstrated the evolutionary dimensional reduction and 2-by-2 global proportionality of phenotypic changes between responses and fluctuations, and between perturbations due to environmental changes or noise and genetic changes, as schematically shown in Fig. 3.
We expect that this evolutionary dimensional reduction and the resulting constraint on phenotypic evolution will be universal in evolved biological systems. In the theoretical argument and cell models, the prerequisites for the emergence of dimensional reduction are only the following two: (i) phenotypes with higher fitness are shaped by complex high-dimensional dynamical systems, which involve stochasticity and (ii) the fraction of such fit states is rare in the state space and in the genetic-rule space. These two features are necessary for our theoretical argument, while most biological systems satisfy them. Intuitively, a high fitness state satisfying macro-micro consistency must be rare, and once it is achieved, its conservation is favored, implying the requirement of robustness.
@@ 6 Universality
The universality of evolutionary dimension reduction, the proportionality between responses to environmental and genetic changes, and the proportionality between genetic variance () and isogenic phenotypic variation () have recently been examined in theoretical models and experiments. These are briefly reviewed here.
@@
Dimensional reduction by evolution has been observed in several models. First, in reaction network models, phenotypic change is mainly constrained along a one-dimensional manifold by different types of fitness for selection (e.g., by selecting the concentration of a component). Second, even when environmental conditions (e.g., the concentration of external nutrients) are not fixed but fluctuate over generations, the constraint is observed within a one-dimensional manifold [6]. Third, by using the gene expression dynamics model and imposing a fitness condition depending on the expression pattern of target genes, dimensional reduction and the relationship between variances are confirmed[8].
@@
Through a combination of data analysis, elastic network models, and theoretical analysis, we have confirmed the relationship between and , as well as the dimensional reduction of protein structure. Changes in estimated structure due to conformation dynamics under noise and genetic mutations are constrained to a common low-dimensional space, typically about five dimensions, implying a strong reduction from the hundreds of residues that give rise to the high-dimensional state space [23]. A theoretical explanation is proposed there, whereas the consequence of few slow modes has been discussed earlier [24], and some model studies also suggest such dimensional reduction [25].
@@
In addition, we have also studied the evolving spin-Hamiltonian system as a simple, abstract model for the evolution of protein dynamics. Here, the spin variable takes up (+1) or down (-1) values whose dynamics is governed by the interaction with other spins, with the energy between the spins and . (If is positive (negative), the two spins tend to align (oppose)). We formulate the above interaction using a Hamiltonian () for energy, with the equilibrium distribution under temperature (). By introducing the fitness determined by a particular configuration of target spins, the matrix is evolved by mutation and selection [26]. The numerical results as well as the statistical-physical calculations show that the evolutionary dimensional reduction and the - relation occur under an intermediate temperature range[27, 28].
(III) Experimental confirmations:
@@
Dimensional reduction was observed in the transcriptome analysis of bacteria E. coli, under different types of stress[2]. Support for equation (1) was obtained by the bacterial evolution under ethanol stress, by using the transcriptome analysis of expression levels of mRNAs[4, 5]. In addition, the transcriptome analysis showed that changes in expression levels of thousands of mRNAs under different antibiotics could be predicted by about seven principal components [9]. The cross-fitness after evolution under different types of antibiotics is shown to be correlated with the response in prior to evolution, as predicted by the theory[8]. Further support for such dimensional reduction comes from Raman analysis of cellular components [29].
@@
The proportionality between genetic and isogenic phenotypic variances across genes is also consistent with measurement of yeast gene expressions. Landry et al.[30] made global measurements of isogenic expression variation and mutation-induced variance. Although they used βmutational varianceβ (i.e., the rate of diffusive spread of each gene expression level upon the introduction of mutations) instead of , a high correlation between the isogenic phenotypic fluctuation () and the mutational variance was observed across all genes in yeast (see also [31]). More recently, the quantitative analysis of the gene expression pattern in Japanese medaka (fish) [32] has suggested - proportionality.
@@
The validity of the proportionality between variances is not limited to gene expressions, but can be applied to any other phenotypic trait. The correlation between the isogenic variance and the rate of evolution was confirmed in the experiment on the evolution of RNA structure[33]. Correlation between genetic and antigenic (βphenotypicβ) changes was noted in influenza evolution[34]. In a series of experiments, Stearns et al.[35] measured the isogenic variance of five life-history traits in Drosophila melanogaster, while also measuring the genetic variance in the same five traits between different genetic lines. Interestingly, they observed proportionality between the isogenic variance of each trait and its genetic variance. Further support for - proportionality can be found in the wing morphology of Drosophila (fly) [36, 37].
(IV) Dimensional reduction in related topics:
@@
The dimensional reduction has been discussed in morphology of shells[38], beaks[39], wing patterns[40] and so forth. It has also been observed in the behavior of microorganisms[41], C. elegans[42, 43], development of C. elegans[44] and in ecological evolution[45]. One possible link to the dimensional reduction will be the sloppy parameter hypothesis[46], which suggests that many parameters used in biological models are irrelevant.
@@
Besides the issue in biophysics, dimension reduction is attracting much attention in learning in the brain [47, 48] and neural networks [49],as well as in machine learning[50]. In deep learning, the possible network configurations that can support a desired output are quite redundant, and reduction to the appropriate output space is achieved. This is similar to our case. The genetic space (e.g., the network configuration) is high-dimensional, so there can be many ways to achieve the appropriate phenotype. Even though the final selected states are constrained in low-dimensional space, the use of high-dimensional space is essential for both machine learning and evolution.